Find a general solution of the system x(t) = Ax(t) for the given matrix A. 12 51 A= -3 - 12
Answer #1

First vector we need to find the eigenvalue and ergen. of A. A= -12 (12 si 23.09 1A-) = 0 = 12- o -3 The charactetic polynomiEigen vector u,= 4+1 t and V2= J 441 This gives us a complex - weilued solution - z = -Bit e 4ti ㅗ -4ti 2,= (Cos 3t i Sin3t)